10. Do the following unit conversion: 1 kilogram = ______ centigram (3)

A) 100 C) 10-5 E) 105

B) 103 D) 103


11. Do the following unit conversion: 3212 feet = ______ kilometer (1 m = 3.28 ft)       (4)

A) 10 C) 88.214 E) 32.12

B) 0.979 D) 17.7


12. Consider a highway with a speed limit of 55 mi/h. Is this speed less than, equal, or greater than 55 km/h?  (1 km = 0.62 miles)  (2)

A) Less than 55 mi/h C) Equal to 55 mi/h

B) Greater than 55 mi/h D) Impossible to say 


13. The acceleration of gravity is approximately 9.81 m/s2 (depending on your location). What is the acceleration of gravity in feet per second squared? ( 1 m = 3.28 ft) (4)

A) 11.37 C) 91.8 E) 32.2

B) 0.981 D) 3.41



1
Expert's answer
2021-06-09T08:04:56-0400

10) 1 kg is equal to "10^5" centigram. Correct answer is (E) "10^5" centigram.

11)

"3212\ ft\cdot\dfrac{1\ m}{3.28\ ft}\cdot\dfrac{1\ km}{1000\ m}=0.979\ km."

Correct answer is (B) 0.979 km.

12)

"55\ \dfrac{km}{h}\cdot\dfrac{0.62\ mi}{km}=34.1\ \dfrac{mi}{h}."

Correct answer is (A) Less than 55 mi/h.

13)

"9.81\ \dfrac{m}{s^2}\cdot\dfrac{3.28\ ft}{1\ m}=32.2\ \dfrac{ft}{s^2}."

Correct answer is (E) "32.2\ \dfrac{ft}{s^2}."